Matt is building a new barn, with length 10 yards more than the width. While determining the footprint of the barn, he measured the diagonal as 50 yards. What will be the dimensions of the barn?
Answer by ikleyn(52864)

.
Answer.  Dimensions are 30 yards by 40 yards.


Classic  (3,4,5)-right angled triangle.


Algebra solution

x^2 + (x+10)^2 = 50^2

x^2 + x^2 + 20x + 100 = 2500

2x^2 + 20x + (100-2500) = 0

2x^2 + 20x - 2400 = 0

x^2 + 10x - 1200 = 0

(x-30)*(x+40) = 0


The only positive root  x= 30  is meaninful, leading to the answer.

Solved.